Circular No. 5605

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1992 No. 2 G. S. Rossano, R. J. Rudy, R. C. Puetter and D. K. Lynch, Space

and Environment Technology Center, The Aerospace Corporation, report: “Observations of Nova Sgr 1992-2 have been obtained using the cooled Aerospace near-infrared scanning germanium spectrometer with a 10′′ × 8′′ oval aperture on the 3-m Shane telescope of the Lick Observatory. Spectroscopic observations were made in the wavelength interval 0.75-1.35 μm at a resolution of 300-550. The unusual spectrum is dominated by neutral oxygen at 844.6 and 1128.7 nm excited by the Ly β flourescence mechanism. Moderate O I at 777.4 nm, 1316.4 nm and other weaker neutral oxygen lines were also detected. Moderately strong Paschen β through ϵ and weak He I 1083.0 nm were also present. The spectrum possesses numerous moderate and weak lines of neutral nitrogen. Weak neutral carbon lines may also be present. Several lines remain unidentified. We do not confirm the presence of the infrared Ca II triplet but instead identify a blend of nitrogen lines in that wavelength interval. The nova was observed on the four nights Aug. 16.2, 17.2, 18.3 and 20.3 UT, a 25-percent decrease in brightness being detected over this time interval. A monochromatic magnitude of 8.9 at 1.25 μm was observed on Aug. 20.3 UT.”

NOVA CYGNI 1992 G. S. Rossano et al. also report that Nova Cyg 1992 was observed in

the same manner as Nova Sgr 1992-2 on Aug. 16.4, 17.4 and 20.4 UT. It was observed to be more typical in appearance, with strong lines of the Paschen series and strong He I 1083.0 nm, along with other lines of He I, He II, O I, [O II], [Ar III], [N I], [S II], [S III] and [P II]. A monochromatic magnitude of 8.4 at 1.25 μm was observed on Aug. 20.4 UT.

Unfiltered CCD magnitudes by W. Liller, Viña del Mar, Chile: Aug. 22.99 UT, 7,91; 24.02, 8.04, 30.98, 8.89, 31.98, 8.77. Comparison star SAO 227635, mag 8.10.

1992 QN Corrigendum. On IAUC 5602 the discoverers should be listed as E. F.

Helin and J. Alu.

1992 September 5 Brian G. Marsden